OpenStudy (anonymous):

on a 40-question test, Peggy answers 34 questions correctly. What percent of the questions does she answer correctly?

OpenStudy (anonymous):

You set up your question like 34/40 = x/100 and then cross multiply and divide to find x

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

You're welcome

OpenStudy (anonymous):

icant seem to get it can i just have the answer please?

OpenStudy (anonymous):

It's really simple. 34/40=x/100 Cross multiply and divide that^ 34(100)=3,400 3,400/100= your answer